Solve by the elimination method
7r - 4s = 24
4r + 7s = 88
What is the solution of the system???


Solve 8x - (6x + 6) =8
The solution is x=??

Answer by mangopeeler07(462)

1. Solve by the elimination method
7r - 4s = 24
4r + 7s = 88
What is the solution of the system???

First of all, you have to make something that you can eliminate. Multiply the second equation by 7/4.

7r - 4s = 24
7r + 49/4s = 154

Now subtract the second equation from the first and get
33/4s=154

Mulitply both sides by 4/33
s=616/33
or 18.666666

Plug that back into the first equation and get

7r - 4(616/33) = 24
or 7r-74.666664=24

Add 74.666664 to both sides and get
7r=98.666667

Divide both sides by 7 and get
r=14.095237
-------------------------------------------------------------
2. Solve 8x - (6x + 6) =8
The solution is x=??

8x-(6x+6)=8
8x-(6x+6)=8x-6x-6

So 8x-6x-6=8

Combine like terms
2x-6=8

Add 6 to both sides
2x=14

Divide both sides by 2
x=7